1 painter can paint a house in 12 hours. Another painter can paint a house in 8 hours. How much time will it take for both painters to paint the house together?
Answer by richwmiller(17219)

1/a+1/b=1/x
1/8+1/12=1/x
1/8+1/12=1/4.8
1/x=1/4.8
x=4.8 days working together
check
1/a+1/b=1/x
1/8+1/12=1/x
12/96+8/96=1/x
20/96=1/x
96/20=x
x=4.8 days working together
ok